RC Meeting, Wednesday, October 06 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  1. Beam was restored yesterday (mid-Swing) and has been very stable since maintenance/beam recovery. Only minor 1.5-hr downtime during today Owl and a 15-min. downtime recently ~8:45 to 9:00

Hall C Status

  1. Oct 05 Day (yesterday):
    • During accelerator maintenance day, Hall C did a full survey for Ellen Becker to investigate the underlying cause of the SHMS frequent interlocks: SHMQ Q1 Interlock. She concluded that the most likely cause is the use of a bad brand of power supply, as the two most recent interlocks had the same symptoms on the faulty power supply. She replaced it with a new brand of power supply. See log-entry SHMS Q1 Interlock Fixed
  2. Oct 05 Swing (yesterday) --> Oct 06 Owl/Day (today)
    • Peaceful PION-LT production data-taking since beam restoration after maintenance day yesterday.
  3. Ciprian Gal will start as Hall C Run Coordinator today.

Current Run Plan

  1. Hall C BCM Calibration has been postponed until further notice from Dave Mack.
  2. Continue taking production data on LH2 target with 45 uA beam at kinematics (SHMS: 8.035 GeV/c, 5.505 deg | HMS: -1.026 GeV/c, 47.07 deg)
    • As of 08:52 today (end of Run 12489, Day Shift), we have accumulated: 2764 pions ( 27.64 % )

Current Issues

  1. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.
  2. Radiation rope set by RADCON is attached to HMS
    • Stephen Kay checked, and the the rope is actually part of a spool attached to the HMS so that when it rotates, the spool extends. ( CLEARED )

Special Requests

  1. Contact Ellen Becker if there is an opportunity to access the hall (~ 4 hours). This is for diagnosing the SHMS power supply. ( CLEARED )
    • Given the recurrent magnet interlocks in the Oct 05 Owl Shift, and that the Accelerator was down for maintenance, Ellen went into the Hall to diagnose the SHMS recurring Interlock issue. As of 13:00 yesterday (Oct 05 Day Shift), Ellen has replaced the faulty power supply with a new brand, and she suggests that the new brand of power supply is used, as it is suspected that the issue is with the previous brand. She found the two most recent SHMS interlocks were due to the same observed symptoms in two low-voltage power supplies of Q1, so she replaced it with another brand. See log entry --> SHMS Q1 Interlock Fixed